In a certain community, it is accepted that 85% of households are two-income households. You



select a random sample of 150 households. What is the probability that more than 130 of them are



two income households?



O 0.1977



O 0.2843



O 0.8023



O 0.7157



O 0.5596

1
Expert's answer
2021-12-23T04:56:05-0500

If "X" is a binomial random variable with mean "\u03bc = np" and variance "\u03c3^2 = npq," then the limiting form of the distribution of "Z=\dfrac{X-np}{\sqrt{npq}}," as "n\to \infin," is the standard normal distribution "n(z;0,1)."

In practice, the approximation is adequate provided that both "np\geq 10" and

"nq\geq 10."

Given "n=150, p=0.85, q=1-p=0.15"


"np=150(0.85)=127.5\geq 10"

"nq=150(0.15)=22.5\geq 10"

The approximation can safely be applied.

Use Normal Distribution with Continuity Correction


"P(X>130)\approx1-P(X<130.5)"

"=1-P(Z<\dfrac{130.5-127.5}{\sqrt{150(0.85)(0.15)}})"

"\approx1-P(Z<0.6860)\approx0.246358"

If we don't use Continuity Correction


"P(X>130)\approx1-P(Z<\dfrac{130-127.5}{\sqrt{150(0.85)(0.15)}})"

"\approx1-P(Z<0.571662)\approx0.2838"